OpenStudy (anonymous):

The first side of a triangle is 8 my shorter than the second side. The third side is 4 times as long as the first side. The perimeter is 26 m. find the length of each side

OpenStudy (dumbcow):

a = b-8 ---> b = a+8 c = 4a a+b+c = 26 using substitution: a + (a+8) + (4a) = 26 solve for "a"